# Fast Estimation of Wasserstein Distances via Regression on Sliced Wasserstein Distances

### v1

@@ 5 Conclusions
Appendix A Proofs
A.1 Proof of Proposition 1
@@
$=\nabla_{\bm{\omega}}\mathbb{E}\left[\bm{\omega}^{\top}\bm{S}_{p}^{(k)}(\mu,\nu)^{\top}\bm{S}_{p}^{(k)}(\mu,\nu)^{\top}\bm{\omega}\right]-2\nabla_{\bm{\omega}}\mathbb{E}\left[\bm{S}_{p}^{(k)}(\mu,\nu)^{\top}\bm{\omega}W_{p}(\mu,\nu)\right]$ (13)
$=\mathbb{E}\left[\nabla_{\bm{\omega}}\bm{\omega}^{\top}\bm{S}_{p}^{(k)}(\mu,\nu)^{\top}\bm{S}_{p}^{(k)}(\mu,\nu)^{\top}\bm{\omega}\right]-2\mathbb{E}\left[\nabla_{\bm{\omega}}\bm{S}_{p}^{(k)}(\mu,\nu)^{\top}\bm{\omega}W_{p}(\mu,\nu)\right]$ (14)
@@ B.1 Gaussian Simulation
Figure 4: Optimal $w^{*}$ and $R^{2}$ in each dimension
@@ B.3 Metric Space Visualization
Data Preprocessing. We follow the same preprocessing pipeline as haviv2024wasserstein .
@@
  • ShapeNetV2 Point Clouds. We use ShapeNetCore.v2 with 15k points per shape. Each shape is normalized to fit inside a unit cube with coordinates in $1,1]^{3}$. We then split each shape into 10k training points and 5k test points, and randomly sample 2,0482{,}048 points from each split.
@@
Wormhole training hyperparameters. We follow the Transformer autoencoder setup of Wormhole with the configuration below:
@@ B.5 RG-Wormhole: Accelerating Wormhole with Regression of Wasserstein
Experimental Settings. We run five experiments to show that RG-Wormhole is much faster than Wormhole with similar effectiveness. First, we measure training time by training both models under the same optimizer and schedule, sweeping batch sizes from 4 to 20 and reporting wall-clock time for training sets of 10, 50, 100, and 200 pairs. Second, we assess encoders by computing $R^{2}$/MSE/MAE between pairwise distances in the learned embedding space and exact Wasserstein. Third, we evaluate decoders by reporting the Wasserstein loss between each input and its reconstruction. Fourth, we examine barycenters by decoding the mean embedding of each class and visualizing results. Fifth, we study interpolation by decoding linear paths between two embeddings and illustrating trajectories. Across all experiments, hyperparameters match Wormhole; the only change in RG-Wormhole is replacing Wasserstein in encoder and decoder losses with the calibrated unconstrained RG. We use 10 samples from the training set to estimate RG coefficients. All experiments use ModelNet40.
@@
Figure 15: ShapeNetV2: RG-Wormhole (constrained model) vs. Wormhole
@@
Figure 16: ShapeNetV2: fast-Wormhole (unconstrained model) vs. Wormhole
Figure 17: ModelNet40: RG-Wormhole vs Wormhole reconstruction experiment
Figure 18: ModelNet40: RG-Wormhole reconstruction experiment
Figure 19: ModelNet40: RG-Wormhole barycenter experiment
Figure 20: ModelNet40: RG-Wormhole barycenter experiment

### v2

@@ 5 Conclusions
Appendix A Details
A.1 Details of Remark 3
@@
$=\nabla_{\bm{\omega}}\mathbb{E}\left[\bm{\omega}^{\top}\bm{S}_{p}^{(k)}(\mu,\nu)\bm{S}_{p}^{(k)}(\mu,\nu)^{\top}\bm{\omega}\right]-2\nabla_{\bm{\omega}}\mathbb{E}\left[\bm{S}_{p}^{(k)}(\mu,\nu)^{\top}\bm{\omega}W_{p}(\mu,\nu)\right]$ (13)
$=\mathbb{E}\left[\nabla_{\bm{\omega}}\bm{\omega}^{\top}\bm{S}_{p}^{(k)}(\mu,\nu)\bm{S}_{p}^{(k)}(\mu,\nu)^{\top}\bm{\omega}\right]-2\mathbb{E}\left[\nabla_{\bm{\omega}}\bm{S}_{p}^{(k)}(\mu,\nu)^{\top}\bm{\omega}W_{p}(\mu,\nu)\right]$ (14)
@@ B.1 Gaussian Simulation
Figure 4: Optimal $w^{*}$ and $R^{2}$ in each dimension.
@@ B.3 Metric Space Visualization
Data Preprocessing. We follow the same preprocessing pipeline as Haviv et al. (2024).
@@
  • ShapeNetV2 Point Clouds. We use ShapeNetCore.v2 with 15k points per shape. Each shape is normalized to fit inside a unit cube with coordinates in $1,1]^{3}$. We then split each shape into 10k training points and 5k test points, and randomly sample 2,0482{,}048 points from each point cloud.
@@
Wormhole training hyperparameters. We follow the Transformer autoencoder setup of Wormhole with the configuration in Table LABEL:tab:config.
@@ B.5 RG-Wormhole: Accelerating Wormhole with Regression of Wasserstein
Experimental Settings. We run five experiments to show that RG-Wormhole is much faster than Wormhole with similar effectiveness. First, we measure training time by training both models under the same optimizer and schedule, sweeping batch sizes from 4 to 20 and reporting wall-clock time for training sets of 10, 50, 100, and 200 pairs. Second, we assess encoders by computing $R^{2}$/MSE/MAE between pairwise distances in the learned embedding space and exact Wasserstein. Third, we evaluate decoders by reporting the Wasserstein loss between each input and its reconstruction. Fourth, we examine barycenters by decoding the mean embedding of each class and visualizing results. Fifth, we study interpolation by decoding linear paths between two embeddings and illustrating trajectories. Across all experiments, hyperparameters match Wormhole; the only change in RG-Wormhole is replacing Wasserstein in encoder and decoder losses with the calibrated unconstrained RG. We use 10 samples from the training set to estimate RG coefficients. Except for embedding experiment which uses ShapeNetV2 dataset, other experiments use ModelNet40 dataset, same as (Haviv et al., 2024).
@@
Figure 15: ShapeNetV2: RG-Wormhole (constrained model) vs. Wormhole.
@@
Figure 16: ShapeNetV2: RG-Wormhole (unconstrained model) vs. Wormhole
Figure 17: ModelNet40: RG-Wormhole vs Wormhole reconstruction experiment.
Figure 18: ModelNet40: RG-Wormhole reconstruction experiment.
Figure 19: ModelNet40: RG-Wormhole barycenter experiment.
Figure 20: ModelNet40: RG-Wormhole barycenter experiment.
B.6 Comparison between RG and classical approximation methods.
Experimental Settings. We conducted direct runtime and accuracy comparisons between our methods and two widely used classical OT accelerations, Sinkhorn and Linear OT, under both single-pair computation and large-scale pairwise computations. For Sinkhorn and Linear OT, no training phase is required. For example, in RG-s, we estimate the optimal regression weight using 5 randomly selected samples ($M=10$ training pairs). The training cost (2.891 seconds (s)) is not included in the inference-time table below, as we focus on inference time only. For Sinkhorn, we conduct a grid search over the entropic regularization hyperparameter in {0.01, 0.05, 0.1} and the number of iterations in {500, 1000}, and report the best-performing configuration for each dataset. We do not exceed 1,000 iterations because the runtime already surpasses that of exact Wasserstein.
For inference runtime comparison, we use point clouds from ShapeNetV2, where each object consists of 2,048 points in 3D, and measure inference time as the number of pairwise computations in seconds increases (see Table 5). These time measurements are relative and may not be precise due to system conditions.
For performance comparison, we evaluate each method’s approximation quality relative to exact Wasserstein using $R^{2}$ in Table 6. Each value is computed over 5,000 test pairs across four datasets: MNIST point clouds, ShapeNetV2, MERFISH cell niches, and scRNA-seq.
Results. From the results above, several observations are clear. First, Sinkhorn performs the worst among the compared methods. Even with 1,000 iterations, it fails to converge on MNIST point clouds and ShapeNetV2 (negative $R^{2}$), while also being slower than both Linear OT and RG-s. However, this may be due to a suboptimal choice of the entropic regularization parameter. In theory, the time complexity of Sinkhorn is $\mathcal{O}((N-M)^{2}n^{2}/\epsilon)$ for $(N-M)^{2}$ pairs. Therefore, Sinkhorn is computationally expensive in this setting. Second, Linear OT provides reasonable accuracy and moderate speed, but its approximation quality is consistently below that of RG-s, and its runtime grows much faster as the number of distributions increases. As discussed, the time complexity of Linear OT is $\mathcal{O}((N-M)n^{2}/\epsilon)$. Third, the RG methods provide more accurate estimates than both Sinkhorn and Linear OT. In addition, RG-s, RG-e, RG-o, and RG-seo are all faster than Linear OT. RG-seo achieves a speed comparable to Linear OT while offering better approximation quality, whereas optimization-based SW variants require substantially more computation time.
Table 5: Comparison in inference runtime among RG and classical approximation methods.
Method 2 Samples (1 pair) 10 Samples (45 pairs) 50 Samples (1,225 pairs) 100 Samples (4,950 pairs) 200 Samples (19,900 pairs)
Wasserstein 0.604s 13.351s 354.823s 1,427.276s 5,030.966s
Sinkhorn (iters=500) 0.441s 10.444s 286.203s 1,154.826s 4,628.393s
Sinkhorn (iters=1,000) 0.482s 21.012s 572.441s 2,308.735s 9,336.621s
Linear OT 0.634s 3.141s 15.473s 31.121s 92.134s
RG-s 0.002s 0.090s 2.043s 8.105s 21.408s
RG-e 0.0005s 0.023s 1.737s 7.060s 18.429s
RG-o 0.013s 0.136s 6.634s 15.310s 54.900s
RG-se 0.002s 0.113s 3.780s 15.170s 39.950s
RG-seo 0.013s 0.147s 7.017s 28.840s 94.200s
Table 6: Comparison in performance among RG and classical approximation methods.
Method MNIST Point Cloud (10,000 test pairs) ShapeNetV2 (10,000 test pairs) MERFISH Cell Niches (10,000 test pairs) scRNA-seq (10,000 test pairs)
Sinkhorn -1.970 -1.549 0.764 0.964
Linear OT 0.697 0.918 0.842 0.984
RG-s 0.925 0.942 0.964 0.989
RG-e 0.910 0.919 0.961 0.991
RG-o 0.767 0.753 0.809 0.923
RG-se 0.933 0.948 0.984 0.996
RG-seo 0.934 0.948 0.969 0.992
B.7 Experiment of Inter-class and Intra-class for RG.
Experimental Settings. We conducted an experiment to test how well RG generalizes across intra-class and inter-class pairs. We trained RG only on intra-class pairs and tested on inter-class pairs, and vice versa, using five scenarios for each setting on ShapeNetV2.
Concretely, for the intra-class $\rightarrow$ inter-class setting, $(A,A)\rightarrow(A,B)$, we ran five experiments. In each experiment, we sampled 10 intra-class training pairs from $(A,A)$, estimated the regression weight, and then constructed a test set by sampling 50 objects from class $A$ and 50 objects from class $B$. This yields 100 distributions in total and therefore 4,950 inter-class test pairs $(A,B)$.
For the inter-class $\rightarrow$ intra-class setting, $(A,B)\rightarrow(A,A)$, we again ran five experiments. In each experiment, we sampled five objects from class $A$ and five objects from class $B$ to form 10 inter-class training pairs $(A,B)$, estimated the regression weight, and then built a test set by sampling 100 objects from class $A$, which yields 4,950 intra-class test pairs $(A,A)$.
Results. We present the results in Figures 2122. Although this is intentionally suboptimal (the regression weight is learned from a restricted subset of pairs), RG still achieves reasonably high $R^{2}$ (e.g., 0.902, 0.846, 0.932, 0.805), compared to about 0.95 when training on the full pair distribution. This shows that RG generalizes well to out-of-sample data, with performance degradation that is present but moderate.
Figure 21: ModelNet40: Intra-class Experiment.
Figure 22: ModelNet40: Inter-class Experiment.
B.8 Experiment of Optimal Weights of RG.
Results. We visualized the estimated coefficients of the RG methods for all applications in Figures 2330. The coefficient patterns vary substantially across datasets, indicating that there is no fixed set of coefficients that universally applies to all cases. This further underscores the need for the proposed regression framework, which captures the dataset-specific uncertainty in the association between Wasserstein and sliced Wasserstein distances (i.e., the meta-distribution) on the full pair distribution.
Figure 23: MNIST Point Cloud: Optimal weight of RG variants (constrained) across different training samples.
Figure 24: MNIST Point Cloud: Optimal weight of RG variants (unconstrained) across different training samples.
Figure 25: ShapeNetV2: Optimal weight of RG variants (constrained) across different training samples.
Figure 26: ShapeNetV2: Optimal weight of RG variants (unconstrained) across different training samples.
Figure 27: MERFISH Cell Niches: Optimal weight of RG variants (constrained) across different training samples.
Figure 28: MERFISH Cell Niches: Optimal weight of RG variants (unconstrained) across different training samples.
Figure 29: scRNA-seq Atlas: Optimal weight of RG variants (constrained) across different training samples.
Figure 30: scRNA-seq Atlas: Optimal weight of RG variants (unconstrained) across different training samples.